# Spaces of subgroups of toral groups

### v1

@@ 10.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)\longrightarrow H^{3}(W;S)$
@@ 10.D. Summary
[…]
@@
[…]
@@
10.E. The view inside $U(2)$ torus normalizer
To start with, we choose the maximal torus $T^{2}$ to consist of the diagonal matrices. The Weyl group $W$ of order 2 acts on this by exchanging the entries. We write $Z$ for scalar matrices, and note that it is the centre, and it is the factor $T$ in the decomposition $U(2)=T\times_{C_{2}}SU(2)$. We write $\tilde{T}$ for $T^{2}\cap SU(2)$ (the matrices $\mathrm{diag}(\lambda,\lambda^{-1})$). This is a maximal torus for $SU(2)$, and $L\tilde{T}$ is the $-1$ eigenspace for the action of $W$. We note that $T^{2}$ is generated by $Z$ and $\tilde{T}$ but $Z\cap\tilde{T}$ is of order 2, so that $T^{2}=Z\times_{C_{2}}\tilde{T}$.
@@ 11.A. Preamble
This example is the normalizer of the maximal torus inside a group of local type $A_{1}\times A_{1}$ (such as $SU(2)\times SU(2)$). In other words $\Lambda_{0}$ is the $A_{1}\times A_{1}$-lattice and taken with the Lie theoretic Weyl group $W=D_{4}\cong C_{2}\times C_{2}$ acting on the lattice. In concrete terms $D_{4}$ acts on $\Lambda_{0}=\tilde{\mathbb{Z}}_{x}\oplus\tilde{\mathbb{Z}}_{y}$ by reflection in the $x$-axis and the $y$-axis (where ~$\tilde{\mathbb{Z}}_{x},\tilde{\mathbb{Z}}_{y}$ consist of integral points on the $x$ and $y$ axes). We have $H^{3}(W;\Lambda_{0})\cong(\mathbb{Z}/2)^{4}$, where the large number of cases arises because of the choice of which element of order 2 in $W$ lifts to an element of order 4, and which element of order 2 in $T^{2}$ is a square in $G$.
@@ 11.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)\longrightarrow H^{3}(W;S)$
@@
Now consider the case when $S$ is finite, and $\Lambda$ has rank 2. The Type 1 lattices are isomorphic to $\Lambda^{0}$ and hence of the form ~$\tilde{\mathbb{Z}}_{x}\oplus\tilde{\mathbb{Z}}_{y}$, with cohomology as above. The behaviour depends on the parities of $m$ and $n$. If both are odd, only the split extension lifts; it lifts uniquely and then splits into four conjugacy classes. If both $m$ and $n$ are even, all classes lift each one in four ways, and each forms a single conjugacy class. If one is even and one is odd, half the classes lift; each lifts in two ways and each of those splits into two conjugacy classes.
The Type 2 lattices are isomorphic to $\mathrm{ind}_{C_{2}}^{W}\tilde{\mathbb{Z}}$, and hence have $H^{1}(W;T^{2}/S)\cong H^{2}(W;\Lambda^{S})=H^{2}(C_{2};\tilde{\mathbb{Z}})=0$ and $H^{2}(W;T^{2}/S)\cong H^{3}(W;\Lambda^{S})=H^{3}(C_{2};\tilde{\mathbb{Z}})=\mathbb{Z}/2$. If $m+n$ and $m-n$ are both even, all classes lift, and otherwise exactly half othem do. Those classes that lift, do so uniquely, to a single conjugacy class.
@@ 11.D. Summary
In each nonsplit case there are 2 subgroups $H(\infty,n)$ when $n$ is even and 0 or 2 subgroups otherwise, and similarly there are 2 subgroups $H(m,\infty)$ when $n$ is even and 00 or $2$ subgroups otherwise depending on $m$. When $m,n$ are finite there are 4 subgroups for certain pairs $(m,n)$ including those with $m$ and $n$ even from Type 1 lattices and an additional 1 from Type 2 from various pairs $m$, $n$ including those where $m+n$ and $m-n$ are even.
@@ 12.B. Lattices and Weyl groups
Type 2: If $m+n$ and $m-n$ are even we have $\Lambda_{2}(m,n)$ which contains $\Lambda_{1}(m,n)$ as a sublattice of index 2. It has basis $\lambda_{1}=((m+n)/2,(m-n)/2)$ and $\lambda_{2}=((m-n)/2,(m+n)/2)$. We have containments
@@ 12.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)\longrightarrow H^{3}(W;S)$
@@ 12.D. Summary
In the split case there are 2 subgroups $H^{s/ns}(\infty,n)$ or $H^{s/ns}(1,\infty)$. When $m,n$ are finite there are 4 subgroups for each pair $(m,n)$ from Type 1 lattices and if $m+n$ and $m-n$ are event there is an additional one from Type 2).
@@ 13.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)\longrightarrow H^{3}(W;S)$
@@ 14.A. Preamble
This example is the normalizer of the maximal torus inside a group of local type $B_{2}$ (such as $Sp(2)$). In other words $\Lambda_{0}$ is the $B_{2}$-lattice and taken with the Lie theoretic Weyl group $W=D_{8}$ acting on the lattice. In concrete terms $D_{8}$ acts on $\Lambda_{0}=\mathbb{Z}\oplus\mathbb{Z}$ generated by reflection in the axes and diagonals and $H^{3}(W;\Lambda_{0})=\mathbb{Z}/2$.
@@ 14.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)\longrightarrow H^{3}(W;S)$
@@
and for $i=1$ the vertical maps are isomorphisms (it is zero for $i=2,3$). We then claim that the map $H^{i}(W;\Lambda_{0})\longrightarrow H^{i}(W;\Lambda_{S})$ for $i=2,3$ is the same as for $i=1$. We thus refer to the behaviour in Section LABEL:sec:secC4subgps.
@@ 14.D. Summary
=$\mathcal{F}=\{H_{1}(m),H^{\prime}_{1}(m)\;|\;m\geq 1\}\cup\{H_{2}(m),H^{\prime}_{2}(m)\;|\;m\geq 1\},$
@@
=$\mathcal{F}=\{H_{1}(2m),H^{\prime}_{1}(2m)\;|\;m\geq 1\}\cup\{H_{2}(m),H^{\prime}_{2}(m)\;|\;m\geq 1\}$
@@ 15.B. Lattices and Weyl groups
Each rank 2 lattice in $\Lambda^{0}=\mathbb{Z}[\omega]$ is generated as a $W$-module by a single element $a+b\omega$ which may be taken to have argument in $[0,\pi/3)$. All these lattices are isomorphic to $\mathbb{Z}[\omega]$ as $W$-modules. The Weyl group is of order 3.
Proof : We choose a nonzero element $a+b\omega$ of $\Lambda$ of minimal length, and we may suppose the argument is in $[0,\pi/3)$. This generates a lattice of equilateral triangles. If $\Lambda\neq\langle a+b\omega,\omega(a+b\omega)\rangle$ then any other element must be closer to a lattice point than $a+b\omega$ is to the origin. This contradicts the choice of $a+b\omega$. The order of the Weyl group follows from the fact that $\det(M)=3$.
@@ 15.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)\longrightarrow H^{3}(W;S)$
@@ 15.D. Summary
:=$\mathcal{F}:=\{a+b\omega\in\mathbb{Z}[\omega]\;|\;\arg(a+b\omega)\in[0,\pi/3)\}$
together with a point for $G$. As a topological space it is the one point compactification of $\mathcal{F}$.
@@ 16.A. Preamble
We will deal with two different toral lattices in this section. Let $W=D_{6}$ and suppose the toral lattice is $\Lambda_{0}=\mathbb{Z}[\omega]^{\prime}$ or $\mathbb{Z}[\omega]^{\prime\prime}$ where the $32\pi/$ 3 rotation $\rho$ acts as multiplication by $\omega$ and the reflection $\tau$ in the $x$ axis acts by complex conjugation (ie reflection in the line through the number $v=1$) in the primed case and reflection in the line through $v=2+\omega$ in the double primed case. In this case $H^{i}(W;T^{2})=H^{i+1}(W;\Lambda_{0})=H^{i+1}(C_{3};\mathbb{Z}[\omega])^{C_{2}}$ The relevant groups are
@@ 16.D. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)\longrightarrow H^{3}(W;S)$

### v2

@@ 10.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
@@ 10.D. Summary
${H(\infty,1)}$${H(\infty,2)}$${H(\infty,3)}\textstyle{\cdots}$${H(\infty,\infty)}\textstyle{\vdots}\textstyle{\vdots}\textstyle{\vdots}\textstyle{\vdots}$${H_{1}^{s/ns}(4,1)}$${H_{1}^{s/ns}(4,2),H_{2}(4,2)}$${H_{1}^{s/ns}(4,3)}\textstyle{\cdots}$${H^{s/ns}(4,\infty)}$${H_{1}^{s/ns}(3,1),H_{2}(3,1)}$${H_{1}^{s/ns}(3,2)}$${H_{1}^{s/ns}(3,3),H_{2}(3,3)}\textstyle{\cdots}$${H^{s/ns}(3,\infty)}$${H_{1}^{s/ns}(2,1)}$${H_{1}^{s/ns}(2,2),H_{2}(2,2)}$${H_{1}^{s/ns}(2,3)}\textstyle{\cdots}$${H^{s/ns}(2,\infty)}$${H_{1}^{s/ns}(1,1),H_{2}(1,1)}$${H_{1}^{s/ns}(1,2)}$${H_{1}^{s/ns}(1,3),H_{2}(1,3)}\textstyle{\cdots}$${H^{s/ns}(1,\infty)}$
@@
${T\times D_{2}}$${T\times D_{4}}$${T\times D_{6}}\textstyle{\cdots}$${T^{2}\rtimes W}\textstyle{\vdots}\textstyle{\vdots}\textstyle{\vdots}\textstyle{\vdots}$${(C_{8}\times_{2}D_{4}),(C_{8}\times_{2}Q_{4})}$${(C_{8}\times_{2}D_{8}),(C_{8}\times_{2}Q_{8})}$${(C_{8}\times_{2}D_{12}),(C_{8}\times_{2}Q_{12})}\textstyle{\cdots}$${(C_{4}\times O(2)),(C_{8}\times_{2}Pin(2))}$${(C_{6}\times_{2}D_{4}),(C_{6}\times_{2}Q_{4}),H_{2}(3,1)}$${(C_{6}\times_{2}D_{8}),(C_{6}\times_{2}Q_{8})}$${(C_{6}\times_{2}D_{12}),(C_{6}\times_{2}Q_{12}),H_{2}(3,3)}\textstyle{\cdots}$${(C_{3}\times O(2)),(C_{6}\times_{2}Pin(2))}$${(C_{4}\times_{2}D_{4}),(C_{4}\times_{2}Q_{4})}$${(C_{4}\times_{2}D_{8}),(C_{4}\times_{2}Q_{8}),H_{2}(2,2)}$${(C_{4}\times_{2}D_{12}),(C_{4}\times_{2}Q_{12})}\textstyle{\cdots}$${(C_{2}\times O(2)),(C_{4}\times_{2}Pin(2))}$${(C_{2}\times_{2}D_{4}),(C_{2}\times_{2}Q_{4}),H_{2}(1,1)}$${(C_{2}\times_{2}D_{8}),(C_{2}\times_{2}Q_{8})}$${(C_{2}\times D_{12}),(C_{2}\times Q_{12}),H_{2}(1,3)}\textstyle{\cdots}$${(C_{1}\times O(2)),(C_{2}\times_{2}Pin(2))}$
@@
10.E. Component structure
We read the component structure off Lemma 10.1 and Lemma 10.2. There is a trivial group over all subgroups on hte right hand vertical (those containing $\tilde{T}$) and all those corresponding to Type 2 lattices. Over all others it is a group of order 2.
10.F. The view inside $U(2)$ torus normalizer
To start with, we choose the maximal torus $T^{2}$ to consist of the diagonal matrices. The Weyl group $W$ of order 2 acts on this by exchanging the entries. We write $Z$ for scalar matrices, and note that it is the centre, and it is the factor $T$ in the decomposition $U(2)=T\times_{C_{2}}SU(2)$. We write $\tilde{T}$ for $T^{2}\cap SU(2)$ (the matrices $\mathrm{diag}(\lambda,\lambda^{-1})$). This is a maximal torus for $SU(2)$, and $L\tilde{T}$ is the 1-1 eigenspace for the action of $W$. We note that $T^{2}$ is generated by $Z$ and $\tilde{T}$ but $Z\cap\tilde{T}$ is of order 2, so that $T^{2}=Z\times_{C_{2}}\tilde{T}$.
@@ 11.A. Preamble
This example is the normalizer of the maximal torus inside a group of local type $A_{1}\times A_{1}$ (such as $SU(2)\times SU(2)$). In other words $\Lambda_{0}$ is the $A_{1}\times A_{1}$-lattice and taken with the Lie theoretic Weyl group $W=D_{4}\cong C_{2}\times C_{2}$ acting on the lattice. In concrete terms $D_{4}$ acts on $\Lambda_{0}=\tilde{\mathbb{Z}}_{x}\oplus\tilde{\mathbb{Z}}_{y}$ by reflection in the $x$-axis and the $y$-axis (where $\tilde{\mathbb{Z}}_{x},\tilde{\mathbb{Z}}_{y}$ consist of integral points on the $x$ and $y$ axes). We have $H^{3}(W;\Lambda_{0})\cong(\mathbb{Z}/2)^{4}$, where the large number of cases arises because of the choice of which element of order 2 in $W$ lifts to an element of order 4, and which element of order 2 in $T^{2}$ is a square in $G$.
@@ 11.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
@@
Now consider the case when $S$ is finite, and $\Lambda$ has rank 2. The Type 1 lattices are isomorphic to $\Lambda^{0}$ and hence of the form $\tilde{\mathbb{Z}}_{x}\oplus\tilde{\mathbb{Z}}_{y}$, with cohomology as above. The behaviour depends on the parities of $m$ and $n$. If both are odd, only the split extension lifts; it lifts uniquely and then splits into four conjugacy classes. If both $m$ and $n$ are even, all classes lift each one in four ways, and each forms a single conjugacy class. If one is even and one is odd, half the classes lift; each lifts in two ways and each of those splits into two conjugacy classes.
The Type 2 lattices are isomorphic to $\mathrm{ind}_{C_{2}}^{W}\tilde{\mathbb{Z}}$, and hence have $H^{1}(W;T^{2}/S)\cong H^{2}(W;\Lambda^{S})=H^{2}(C_{2};\tilde{\mathbb{Z}})=0$ and $H^{2}(W;T^{2}/S)\cong H^{3}(W;\Lambda^{S})=H^{3}(C_{2};\tilde{\mathbb{Z}})=\mathbb{Z}/2$. If $m+n$ and mnm-n are both even, all classes lift, and otherwise exactly half othem do. Those classes that lift, do so uniquely, to a single conjugacy class.
@@ 11.D. Summary
In each nonsplit case there are 2 subgroups $H(\infty,n)$ when $n$ is even and 0 or 2 subgroups otherwise, and similarly there are 2 subgroups $H(m,\infty)$ when $n$ is even and 0 or $2$ subgroups otherwise depending on $m$. When $m,n$ are finite there are 4 subgroups for certain pairs $(m,n)$ including those with $m$ and $n$ even from Type 1 lattices and an additional 1 from Type 2 from various pairs $m$, $n$ including those where $m+n$ and mnm-n are even.
@@ 12.B. Lattices and Weyl groups
Type 2: If $m+n$ and mnm-n are even we have $\Lambda_{2}(m,n)$ which contains $\Lambda_{1}(m,n)$ as a sublattice of index 2. It has basis $\lambda_{1}=((m+n)/2,(m-n)/2)$ and $\lambda_{2}=((m-n)/2,(m+n)/2)$. We have containments
@@ 12.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
@@ 12.D. Summary
In the split case there are 2 subgroups $H^{s/ns}(\infty,n)$ or $H^{s/ns}(1,\infty)$. When $m,n$ are finite there are 4 subgroups for each pair $(m,n)$ from Type 1 lattices and if $m+n$ and mnm-n are event there is an additional one from Type 2).
@@ 13.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
@@ 14.A. Preamble
This example is the normalizer of the maximal torus inside a group of local type $B_{2}$ (such as $Sp(2)$). In other words $\Lambda_{0}$ is the $B_{2}$-lattice and taken with the Lie theoretic Weyl group $W=D_{8}$ acting on the lattice. In concrete terms $D_{8}$ acts on $\Lambda_{0}=\mathbb{Z}\oplus\mathbb{Z}$ generated by reflection in the axes and diagonals and $H^{2}(W;\Lambda_{0})=\mathbb{Z}/2j,H^{3}(W;\Lambda_{0})=(\mathbb{Z}/2)^{2}$. We will focus the analysis on the split extension. The case of the normalizer of the maximal torus in $Sp(2)$ will be treated elsewhere.
@@ 14.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
@@
and for $i=1$ the vertical maps are isomorphisms (it is zero for $i=2,3$). We then claim that the map $H^{i}(W;\Lambda_{0})\longrightarrow H^{i}(W;\Lambda_{S})$ for $i=2,3$ is the same as for $i=1$. We thus refer to the behaviour in Section 13.
@@ 14.D. Summary
$\mathcal{F}=\{H_{1}(m),H^{\prime}_{1}(m)\;|\;m\geq 1\}\cup\{H_{2}(m),H^{\prime}_{2}(m)\;|\;m\geq 1\},$
@@
$\mathcal{F}=\{H_{1}(2m),H^{\prime}_{1}(2m)\;|\;m\geq 1\}\cup\{H_{2}(m),H^{\prime}_{2}(m)\;|\;m\geq 1\}$
@@ 15.B. Lattices and Weyl groups
Each rank 2 lattice in $\Lambda^{0}=\mathbb{Z}[\omega]$ is generated as a $W$-module by a single element $a+b\omega$ which may be taken to have argument in $0,\pi/3)$. All these lattices are isomorphic to $\mathbb{Z}[\omega]$ as $W$-modules. The Weyl group is of order 3.
Proof : We choose a nonzero element $a+b\omega$ of $\Lambda$ of minimal length, and we may suppose the argument is in $0,\pi/3)$. This generates a lattice of equilateral triangles. If $\Lambda\neq\langle a+b\omega,\omega(a+b\omega)\rangle$ then any other element must be closer to a lattice point than $a+b\omega$ is to the origin. This contradicts the choice of $a+b\omega$. The order of the Weyl group follows from the fact that $\det(M)=3$.
@@ 15.C. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$
@@ 15.D. Summary
$\mathcal{F}:=\{a+b\omega\in\mathbb{Z}[\omega]\;|\;\arg(a+b\omega)\in[0,\pi/3)\}$
together with a point for $G$. As a topological space it is the one point compactification of $\mathcal{F}$. The component structure has a group of order 3 over points of $\mathcal{F}$ and the trivial group over $G$.
@@ 16.A. Preamble
We will deal with two different toral lattices in this section. Let $W=D_{6}$ and suppose the toral lattice is $\Lambda_{0}=\mathbb{Z}[\omega]^{\prime}$ or $\mathbb{Z}[\omega]^{\prime\prime}$ where the $2\pi/3$ rotation $\rho$ acts as multiplication by $\omega$ and the reflection $\tau$ in the $x$ axis acts by complex conjugation (ie reflection in the line through the number $v=1$) in the primed case and reflection in the line through $v=2+\omega$ in the double primed case. In this case $H^{i}(W;T^{2})=H^{i+1}(W;\Lambda_{0})=H^{i+1}(C_{3};\mathbb{Z}[\omega])^{C_{2}}$ . The relevant groups are
@@ 16.D. Sections and conjugacy
$H^{1}(W;T^{2})\stackrel{{\scriptstyle a^{1}}}{{\longrightarrow}}H^{1}(W;T^{2}/S)\longrightarrow H^{2}(W;S)\stackrel{{\scriptstyle i^{S}_{*}}}{{\longrightarrow}}H^{2}(W;T^{2})\stackrel{{\scriptstyle a^{2}}}{{\longrightarrow}}H^{2}(W;T^{2}/S)$